Y"+6y'+9y=-xe^4x by variation of parameters


1
Expert's answer
2021-05-24T00:52:03-0400
"y''+6y'+9y=-xe^{4x}"

The corresponding homogeneous equation is


"y''+6y'+9y=0"

The auxiliary (characteristic) equation is given by 


"r^2+6r+9=0"

"(r+3)^2=0"

There is one repeated real root "r=-3."

The general solution of the homogeneous equation is


"y_c=C_1e^{-3x}+C_2xe^{-3x}"

We have

"y_1=e^{-3x}, y_2=xe^{-3x}, g(x)=-xe^{4x}"

The Wronskian of these two functions is


"W(y_1,y_2)=\begin{vmatrix}
 y_1 & y_2 \\
 y_1' & y_2'
\end{vmatrix}=\begin{vmatrix}
 e^{-3x} & xe^{-3x} \\
 -3e^{-3x} & (1-3x)e^{-3x}
\end{vmatrix}"

"=e^{-3x}(1-3x)e^{-3x}-xe^{-3x}(-3)e^{-3x}"

"=e^{-6x}(1-3x+3x)=e^{-6x}"



"W_1=\begin{vmatrix}
 0 & y_2 \\
 g(x) & y_2'
\end{vmatrix}=\begin{vmatrix}
 0 & xe^{-3x} \\
 -xe^{4x} & (1-3x)e^{-3x}
\end{vmatrix}=x^2e^x"


"W_2=\begin{vmatrix}
 y_1 & 0 \\
 y_1' & g(x)
\end{vmatrix}=\begin{vmatrix}
 e^{-3x} & 0 \\
 -3e^{-3x} & -xe^{4x}
\end{vmatrix}=xe^x"

"u_1=\int\dfrac{W_1}{W}dx=\int\dfrac{x^2e^x}{e^{-6x}}dx=\int x^2e^{7x}dx"

"=\dfrac{1}{7}x^2e^{7x}-\dfrac{2}{7}\int xe^{7x}dx"

"=\dfrac{1}{7}x^2e^{7x}-\dfrac{2}{49}xe^{7x}+\dfrac{2}{49}\int e^{7x}dx"

"=\dfrac{1}{7}x^2e^{7x}-\dfrac{2}{49}xe^{7x}+\dfrac{2}{343}e^{7x}"


"u_2=\int\dfrac{W_2}{W}dx=\int\dfrac{xe^x}{e^{-6x}}dx=\int xe^{7x}dx"

"=\dfrac{1}{7}xe^{7x}-\dfrac{1}{7}\int e^{7x}dx"

"=\dfrac{1}{7}xe^{7x}-\dfrac{1}{49}e^{7x}"

"y_p=u_1y_1+u_2y_2"

"=(\dfrac{1}{7}x^2e^{7x}-\dfrac{2}{49}xe^{7x}+\dfrac{2}{343}e^{7x})(e^{-3x})+"

"+(\dfrac{1}{7}xe^{7x}-\dfrac{1}{49}e^{7x})(xe^{-3x})"

"=(\dfrac{2}{7}x^2-\dfrac{3}{49}x+\dfrac{2}{343})e^{4x}"

The general solution of the homogeneous equation is


"y=y_c+y_p"


"=C_1e^{-3x}+C_2xe^{-3x}+\big(\dfrac{2}{7}x^2-\dfrac{3}{49}x+\dfrac{2}{343}\big)e^{4x}"